a pair of shoes costs 52.99 and the state sales tax is 8%. use the formula C=p+rp to find the total cost of the shoes, where C is the total cost,p is the price, and ris the sales tax rate.
Answer by jessica43(140)

The pair of shoes costs $52.99, so p = 52.99.
The state sales tax is 8%, which can be changed to decimal form by dividing by 100: 8/100 = 0.08. So r = 0.08.
Now you can plug both of these values into the formula to find the total cost of the shoes (c):
c = p + r*p
c = 52.99 + (0.08)*(52.99)
c = 52.99 + 4.2392
c = 57.2292
So the total cost of the shoes is $57.23.